OpenStudy (anonymous):

Which line is a graph of the equation: 2x + 5y = 10? VHS_ALG_S1_07_L203_L303_LQ4_graphi.gif A. line a B. line b C. line c D. line d

OpenStudy (anonymous):

@iambatman

OpenStudy (anonymous):

@satellite73 plz help

OpenStudy (anonymous):

Oh wow, now that's a messy looking graph :P, alright so you have the line: 2x + 5y = 10, do you know y=mx+b form? It's is the equation of a straight line.

OpenStudy (anonymous):

sorry no

OpenStudy (anonymous):

Well for these sorts of questions, one of the best ways to go onto see how the graph is by putting it into the form of y = mx+b where m is the slope, and b is the y - intercept (where the line crosses the y - axis)

OpenStudy (anonymous):

So can you try solving for y?

OpenStudy (anonymous):

you have to say x=0 right?

OpenStudy (anonymous):

Nope, not this time, well you could find the intercepts, but I think it will be more helpful if you solve for y as it is and then graph it out.

OpenStudy (anonymous):

I don't get how you do that

OpenStudy (anonymous):

\[2x+5y=10\] try your best to "isolate y" meaning get it on its own, use what I told you earlier, using opposite operations and if you do something to one side you must also do it to the other.

OpenStudy (anonymous):

so I minus 2 - 10?

OpenStudy (anonymous):

subtract 2x from both sides

OpenStudy (anonymous):

\[2x\color{red}{-2x}+5y=10\color{red}{-2x}\]

OpenStudy (anonymous):

then I divide my answer by 5y?

OpenStudy (anonymous):

Well we are solving for y, so just the 5 :)

OpenStudy (anonymous):

And then it will be in y = mx+b form (equation of a straight line)

OpenStudy (anonymous):

y=1.6

OpenStudy (anonymous):

\[5y = -2x+10 \implies y = \frac{ -2x+10 }{ 5 } \implies y = \frac{ -2 }{ 5 }x+2\] so it will look sort of like this

OpenStudy (anonymous):

because the x tells us how far along we should go, and y tells us how far up, I hope all of this makes sense, if not just ask!

OpenStudy (anonymous):

so the answer is D?

OpenStudy (anonymous):

Not quite

OpenStudy (anonymous):

wait sorry no it would be B, right?

OpenStudy (anonymous):

Yup!

OpenStudy (anonymous):

ok one more plz I will start a new one